38) DESIRED: 75 mg P.O. AVAILABLE: 15 mg per 2 cc GIVE: OZ
Answer #1

ANSWER.

* Calculation of amount(oz) of drug required.

     Formula

        Volume(ml) = Desired dose(D) / Dose in hand (H) * Quantity(Q).

Here

      Desired dose(D) = 75 mg.

      Dose in hand(H) = 15 mg.

            Quantity(Q) = 2 ml (1cc = 1 ml)

As per above formula

           Volume(ml) of drug required = 75/15 * 2

                                                         = 5 * 2

                                                         = 10 ml.

                               1 ml = 0.03 oz

            So

               Volume of drug required in oz = 10 * 0.03

                                                                = 0.3 oz.